The grade appeals process at a university requires that a jury be structured by selecting six individuals randomly from a proof of seven students and six faculty. A) what is the probability of selecting a jury of all faculty? B) what is the probability of selecting a jury of three students and three faculty ?


1
Expert's answer
2021-02-28T07:10:14-0500

Out of "7" students and "6" faculty , a jury of six individuals can be select in "^{13}C_6" ways.

Therefore sample space "S" contains "^ {13}C_6" number of elements.

(A) There are "6" faculty. So out "6" faculty a jury of 6 persons can be select in "^6C_6" ways.

Let "A" be en event such that selecting a jury of all faculty.

Then "P(A)=\frac{n(A)}{n(S)}=\frac{^6C_6}{^{13}C_6}=\frac{1}{1716}"

(B) Let "B" be an event such that selecting a jury of "3" students and "3" faculty.

Now a jury of "3" students and "3" faculty out of "7" students and "6" faculty can be select in "(^7C_3\u00d7{ ^6C_3})" ways.

Then "P(B)=\frac{n(B)}{n(S)}=\frac{(^7C_3×^6C_3)}{^{13}C_6}=\frac{700}{1716}=\frac{175}{429}"